A 24-cm diameter piston sits atop a 54000 cm3 canister of 1 mol of nitrogen at 1 atm. Placing a 100kg object on top of the piston compresses the gas isothermally. How far has the piston moved?
Answer #1

here,

for the isothermal process

let the piston is moved by a distance x

P * V = constant

P1 * V1 = P2 * V2

1.01 *10^5 * 54000 = (1.01 *10^5 + 100 * 9.8/(pi * 0.12^2)) * (54000 - pi * 12^2 * x)

solving for x

x = 21.1 cm

the distance moved by the piston is 21.1 cm